# Stylized facts in Web3

### v2

@@ 1. Introduction
The financial landscape has undergone a significant transformation with cryptocurrencies, spearheaded by Bitcoin (BTC) [Nak08]. The growing integration of cryptocurrencies into investor portfolios and the financial industry’s increasing interest, evidenced by recently created Bitcoin-based ETFs, underscore their rising prominence. However, Bitcoin represents merely the tip of the iceberg in digital finance.
The introduction of the Ethereum network [But13] marked a pivotal moment, enabling the creation of an ecosystem of decentralized applications now collectively known as Web3, with Decentralized Finance (DeFi) as its most successful example [HRS21]. This development has expanded the cryptocurrency offering beyond simple Bitcoin clones to essential components of a new financial paradigm.
In this evolving landscape, the term "cryptocurrency" has become limiting. "Tokens" more accurately describe the diverse array of digital assets. For instance, Ether (ETH) serves as the native token of the Ethereum network, facilitating smart contract execution. Similarly, liquidity provider tokens like CRV allow holders to earn fees for liquidity on decentralized exchanges such as Curve while also being tradable assets. The rapid expansion of DeFi over the past five years has given rise to applications addressing a broad spectrum of financial needs, as detailed in [HRS21].
@@
This paper presents a comprehensive statistical analysis of various tokens within the Web3 ecosystem. We report a series of stylized facts and contrast these with well-documented stylized facts in traditional equity markets [MS99, Con01, BP03, ACA11]. Our analysis aims to characterize the main tokens in circulation thoroughly, potentially guiding future models and contributing to Web3’s further development.
@@ 2.1. Blockchain Technology
Blockchain technology is the cornerstone of Web3 systems, comprising a chain of information blocks interconnected and secured through cryptographic keys [LT22]. This distributed ledger technology ensures decentralized storage and transparency across a global network of computers. While initially serving primarily as a transaction record for Bitcoin [Nak08], blockchain technology has evolved to support more complex functionalities, particularly with platforms like Ethereum.
A primary challenge in decentralized systems is ensuring transaction uniqueness and preventing double-spending. Consensus mechanisms such as "proof-of-work" (PoW) or "proof-of-stake" (PoS) [TS19] address this by requiring network participants (miners or validators) to validate and agree on the ledger’s state.
@@ 3.1. Data Sources and Selection Criteria
Our data sources include Centralized Exchanges (CEX), specifically token prices from Binance and Coinbase, two of the largest and most liquid cryptocurrency exchanges. We also gathered data from Decentralized Exchanges (DEX), focusing on Ethereum (ETH) prices from Uniswap, a leading DEX. Here, we considered 5 basis point (bp) and 30 bp liquidity pools (L1 and L2), examining ETH pairs with major stablecoins: USDC, USDT, and DAI. For comparison with traditional financial markets, we included SPY ETF prices as a proxy for the S&P 500 index and the EUR to USD exchange rate.
@@ 3.3. What Do We Model ?
We choose log-returns because it allows us to benchmark our results against traditional finance assets in the literature, which normally use log-returns. However, it’s important to note that the choice of log-returns has potential drawbacks, particularly for Web3 tokens, given their large volatility. As discussed in [BP03], log-returns can potentially create spurious effects and represent a mathematical abstraction that cannot be directly traded.
@@ 4. Stylized Facts
Stylized facts are well-established empirical observations consistently found in financial market data [MS99, Con01, BP03, ACA11]. They provide a qualitative framework of accepted characteristics for asset and market models, emphasizing the importance of detailed empirical analysis before proposing theoretical models. This approach, akin to experimental sciences like physics, has been fundamental in developing econophysics [MS99].
While several studies have documented stylized facts for cryptocurrencies, particularly Bitcoin and Ethereum (e.g. [ZCCN19, TAHA24]), our study expands this analysis to a broader, curated list of tokens representing various Web3 applications (Figure 1). Notably, we compare Ethereum (ETH) traded on both decentralized (DEX) and centralized exchanges (CEX), offering novel insights into the evolving crypto market structure.
@@ 4.1. Fat-tails and Aggregated Normality
One of the most unequivocal stylized facts in finance is the presence of fat-tails in the unconditional probability distribution of returns for time intervals typically less than a few days. This phenomenon implies that rare events are more likely to occur than what a Gaussian distribution would predict, hence the terms "fat-tails" or "heavy-tails" [ACA11]. It has also been observed that this probability distribution gradually approaches a Gaussian distribution (i.e., tails become less fat) for returns over longer periods, typically weeks or months. We refer to this convergence towards a Gaussian distribution as aggregated normality. In this section, we document these well-established stylized facts for Web3 tokens.
@@ 4.1.1. Probability Density Function Analysis
The figure clearly shows that one-hour returns exhibit fat-tails which are well fit to a power-law ($\sim|x|^{-\alpha-1}$) with $\alpha\sim 2.4$, in agreement with the financial literature [ACA11]. The one-day returns are closer to a standard normal at the center, and the tails have become asymmetric and closer to exponential ($\sim\exp{(-\eta|x|)}$) with $\eta\sim 1.7$, close to the $\sqrt{2}$ predicted by [SPY04] for stocks. Finally, at one month, the returns are well approximated by a standard normal probability distribution (red line).
@@
In Appendix Appendix 1. Tail Behavior Analysis, we analyze the cumulative distribution function (CDF) and fit the tails of one-hour returns to a power law ($\text{CDF}(x)\sim|x|^{-\alpha}$) for returns exceeding two standard deviations. We find that the power coefficient $\alpha>2$, typically between 2 and 3 with an average of 2.6 (Table 1), aligns with traditional finance observations [ACA11].
@@ 4.1.2. Jarque-Bera Test for Normality
$\text{JB}(X)=\frac{n}{6}\left(s^{2}(X)+\frac{1}{4}(k(X)-3)^{2}\right)$
@@ 4.2. Autocorrelation of Returns
Analysis of Figure 5 reveals that while most lags in the ACF are insignificant, validating the stylized fact for our tokens, significant autocorrelations are observed within the first hour and at 24 hours. Notably, Mainnet and Arbitrum Uniswap trading pools (30bp) show significant positive ACF values in the first hour, whereas Binance ETH aligns with SPY and EUR, showing insignificant first-hour ACF. A significant negative autocorrelation is observed at 24 hours, repeating less pronouncedly at 48 hours, echoing findings by [HS08]. These observations suggest distinct behaviors between DEX and CEX environments, as well as potential cyclical patterns in crypto trading.
@@ 4.3. Volatility Clustering
According to [ACA11], typical power-law exponents (slope of a log-log fit) for stocks range between -0.1 and -0.4. Our analysis yields values between -0.15 for AAVE and -0.35 for GNO or -0.34 for SHIB. These values fall well within the expected range for stock markets, further corroborating the stylized fact of volatility clustering in crypto assets.
@@ 4.4. Leverage Effect
Figure 8 illustrates the correlation (Eq. 1) between past returns and future absolute returns on the negative axis and the correlation between past absolute returns and future returns on the positive axis for ETH (CEX and DEX) and traditional assets over the same historical period. The correlation is multiplied by the square root of the number of hours, consistent with other correlation figures in this study.
Figure 8. Cross Correlation (Eq. 1) between absolute return and return where these values are multiplied by the square root of the number of hours. Red dashed horizontal lines show ±$\pm 3$ SE (99%) error bars.
@@
It is well established that stock indices exhibit the leverage effect [BP03], while currencies do not [Zum09]. Based on these literature findings and our results in Figure 8, we can confidently state that ETH shows the leverage effect, albeit with lower significance than other stylized facts in this study. However, we are less certain about generalizing the leverage effects for other tokens.
@@
Figure 9. Cross-correlation (Eq. 1) for all Coinbase tokens in Table 1 with error bars (red dashed horizontal lines). Most $L(k)$ values are within the error bars for $k>0$.
To condense this information, Figure 10 plots the average $L(k)$ for $k<0$ against the average $L(k)$ for $k>0$. The average for $k<0$ (labeled "R->|R|") measures the average strength of the relation between current and future absolute returns (a proxy for volatility). The leverage effect expects this relation to be negative: a large price drop (negative return) is followed by periods of increased volatility (absolute returns). It also expects the average $L(k)$ for $k>0$ (labeled "|R|->R") to be close to zero, indicating that current absolute returns do not forecast future returns.
@@ 4.5. Time Reversal Asymmetry
Time reversal asymmetry (TRA) in financial time series, introduced by [Zum09], is a concept linked to entropy and time directionality. It suggests that financial time series exhibit different statistical features when viewed in reverse chronological order. This phenomenon can be illustrated by the analogy of observing broken glass: while one can easily deduce that glass fragments originated from an intact cup, it’s nearly impossible to envision the original cup when presented only with its fragments.
@@
To quantify TRA, we employ a correlation measure introduced in [OEER20]:
$C(k)\equiv\frac{\mathbb{E}\left[(\lvert R_{t}\rvert-\mu_{\lvert R\rvert})(s^{(D)}{t+k}-\mu{s^{(D)}})\right]}{\sigma_{\lvert R\rvert}\sigma_{s^{(D)}}}$ (2)
@@
An increasing function of $\Delta(N)$ confirms the presence of TRA [OEER20].
Figure 11 illustrates $\Delta(N)$ for selected assets over the same historical period. The consistently increasing lines for all assets, including both crypto and traditional financial instruments (SPY and EUR), validate the presence of TRA, aligning with previous findings in traditional finance [Zum09, OEER20].
Figure 11. Cumulative difference function $\Delta(N)$ of Eq. 3 for ETH in different exchanges as well as SPY and EUR.
@@ 4.6. Correlation Matrix and Market Factors
Figure 14 reveals striking differences between the token and stock markets. In the token universe, the top eigenvalue accounts for 56% of the total variance, whereas in the stock samples, it explains only about 26% (reported conservatively as the average minus three standard errors over the 500 randomly selected stock portfolios). Importantly, the first eigenvector is positive for all tokens in our sample, consistent with its interpretation as the "market" factor, as discussed in [PGR+00, Con01, BP03].
@@
These findings suggest that the token universe is driven by a smaller set of eigenvalues/factors than the stock market while still adhering to the stylized fact of factor-driven markets observed in stocks [PGR+00, Con01, BP03]. This conclusion has important implications for modeling and risk control of token portfolios. Specifically, it suggests that a few factor models—likely fewer than those required for stocks—may be sufficient to effectively model the cross-section of tokens.

### v3

@@ Abstract.
Blockchains, Cryptocurrency, Web3, Decentralized Finance, Econophysics, Stylized Facts
Key words and phrases:
@@ 1. Introduction
The financial landscape has undergone a significant transformation with cryptocurrencies, spearheaded by Bitcoin (BTC) [12]. The growing integration of cryptocurrencies into investor portfolios and the financial industry’s increasing interest, evidenced by recently created Bitcoin-based ETFs, underscore their rising prominence. However, Bitcoin represents merely the tip of the iceberg in digital finance.
The introduction of the Ethereum network [4] marked a pivotal moment, enabling the creation of an ecosystem of decentralized applications now collectively known as Web3, with Decentralized Finance (DeFi) as its most successful example [6]. This development has expanded the cryptocurrency offering beyond simple Bitcoin clones to essential components of a new financial paradigm.
In this evolving landscape, the term "cryptocurrency" has become limiting. "Tokens" more accurately describe the diverse array of digital assets. For instance, Ether (ETH) serves as the native token of the Ethereum network, facilitating smart contract execution. Similarly, liquidity provider tokens like CRV allow holders to earn fees for liquidity on decentralized exchanges such as Curve while also being tradable assets. The rapid expansion of DeFi over the past five years has given rise to applications addressing a broad spectrum of financial needs, as detailed in [6].
@@
This paper presents a comprehensive statistical analysis of various tokens within the Web3 ecosystem. We report a series of stylized facts and contrast these with well-documented stylized facts in traditional equity markets [9, 5, 3, 2]. Our analysis aims to characterize the main tokens in circulation thoroughly, potentially guiding future models and contributing to Web3’s further development.
@@ 2.1. Blockchain Technology
Blockchain technology is the cornerstone of Web3 systems, comprising a chain of information blocks interconnected and secured through cryptographic keys [8]. This distributed ledger technology ensures decentralized storage and transparency across a global network of computers. While initially serving primarily as a transaction record for Bitcoin [12], blockchain technology has evolved to support more complex functionalities, particularly with platforms like Ethereum.
A primary challenge in decentralized systems is ensuring transaction uniqueness and preventing double-spending. Consensus mechanisms such as "proof-of-work" (PoW) or "proof-of-stake" (PoS) [19] address this by requiring network participants (miners or validators) to validate and agree on the ledger’s state.
@@ 3.1. Data Sources and Selection Criteria
Our data sources include Centralized Exchanges (CEX), specifically token prices from Binance and Coinbase  [^2] Code to download data: https://github.com/silvaac/token_style, two of the largest and most liquid cryptocurrency exchanges. We also gathered data from Decentralized Exchanges (DEX), focusing on Ethereum (ETH) prices from Uniswap, a leading DEX. Here, we considered 5 basis point (bp) and 30 bp liquidity pools (L1 and L2), examining ETH pairs with major stablecoins: USDC, USDT, and DAI. For comparison with traditional financial markets, we included SPY ETF prices as a proxy for the S&P 500 index and the EUR to USD exchange rate.
@@ 3.3. What Do We Model ?
We choose log-returns because it allows us to benchmark our results against traditional finance assets in the literature, which normally use log-returns. However, it’s important to note that the choice of log-returns has potential drawbacks, particularly for Web3 tokens, given their large volatility. As discussed in [3], log-returns can potentially create spurious effects and represent a mathematical abstraction that cannot be directly traded.
@@ 4. Stylized Facts
Stylized facts are well-established empirical observations consistently found in financial market data [9, 5, 3, 2]. They provide a qualitative framework of accepted characteristics for asset and market models, emphasizing the importance of detailed empirical analysis before proposing theoretical models. This approach, akin to experimental sciences like physics, has been fundamental in developing econophysics [9].
While several studies have documented stylized facts for cryptocurrencies, particularly Bitcoin and Ethereum (e.g. [20, 17]), our study expands this analysis to a broader, curated list of tokens representing various Web3 applications (Figure 1). Notably, we compare Ethereum (ETH) traded on both decentralized (DEX) and centralized exchanges (CEX), offering novel insights into the evolving crypto market structure.
@@ 4.1. Fat-tails and Aggregated Normality
One of the most unequivocal stylized facts in finance is the presence of fat-tails in the unconditional probability distribution of returns for time intervals typically less than a few days. This phenomenon implies that rare events are more likely to occur than what a Gaussian distribution would predict, hence the terms "fat-tails" or "heavy-tails" [2]. It has also been observed that this probability distribution gradually approaches a Gaussian distribution (i.e., tails become less fat) for returns over longer periods, typically weeks or months. We refer to this convergence towards a Gaussian distribution as aggregated normality. In this section, we document these well-established stylized facts for Web3 tokens.
@@ 4.1.1. Probability Density Function Analysis
The figure clearly shows that one-hour returns exhibit fat-tails which are well fit to a power-law ($\sim|x|^{-\alpha-1}$) with $\alpha\sim 2.4$, in agreement with the financial literature [2]. The one-day returns are closer to a standard normal at the center, and the tails have become asymmetric and closer to exponential ($\sim\exp{(-\eta|x|)}$) with $\eta\sim 1.7$, close to the $\sqrt{2}$ predicted by [16] for stocks. Finally, at one month, the returns are well approximated by a standard normal probability distribution (red line).
@@
In Appendix Appendix 1. Tail Behavior Analysis, we analyze the cumulative distribution function (CDF) and fit the tails of one-hour returns to a power law ($\text{CDF}(x)\sim|x|^{-\alpha}$) for returns exceeding two standard deviations. We find that the power coefficient $\alpha>2$, typically between 2 and 3 with an average of 2.6 (Table 1), aligns with traditional finance observations [2].
@@ 4.1.2. Jarque-Bera Test for Normality
$\text{JB}(X)=\frac{n}{6}\left(s^{2}(X)+\frac{1}{4}(k(X)-3)^{2}\right)$ (1)
@@ 4.2. Autocorrelation of Returns
Analysis of Figure 5 reveals that while most lags in the ACF are insignificant, validating the stylized fact for our tokens, significant autocorrelations are observed within the first hour and at 24 hours. Notably, Mainnet and Arbitrum Uniswap trading pools (30bp) show significant positive ACF values in the first hour, whereas Binance ETH aligns with SPY and EUR, showing insignificant first-hour ACF. A significant negative autocorrelation is observed at 24 hours, repeating less pronouncedly at 48 hours, echoing findings by [7]. These observations suggest distinct behaviors between DEX and CEX environments, as well as potential cyclical patterns in crypto trading.
@@ 4.3. Volatility Clustering
According to [2], typical power-law exponents (slope of a log-log fit) for stocks range between -0.1 and -0.4. Our analysis yields values between -0.15 for AAVE and -0.35 for GNO or -0.34 for SHIB. These values fall well within the expected range for stock markets, further corroborating the stylized fact of volatility clustering in crypto assets.
@@ 4.4. Leverage Effect
Figure 8 illustrates the correlation (Equation (2)) between past returns and future absolute returns on the negative axis and the correlation between past absolute returns and future returns on the positive axis for ETH (CEX and DEX) and traditional assets over the same historical period. The correlation is multiplied by the square root of the number of hours, consistent with other correlation figures in this study.
Figure 8. Cross Correlation (Equation (2)) between absolute return and return where these values are multiplied by the square root of the number of hours. Red dashed horizontal lines show ±$\pm 3$ SE (99%) error bars.
@@
It is well established that stock indices exhibit the leverage effect [3], while currencies do not [21]. Based on these literature findings and our results in Figure 8, we can confidently state that ETH shows the leverage effect, albeit with lower significance than other stylized facts in this study. However, we are less certain about generalizing the leverage effects for other tokens.
@@
Figure 9. Cross-correlation (Equation (2)) for all Coinbase tokens in Table 1 with error bars (red dashed horizontal lines). Most $L(k)$ values are within the error bars for $k>0$.
To condense this information, Figure 10 plots the average $L(k)$ for $k<0$ against the average $L(k)$ for $k>0$. The average for $k<0$ (labeled "$R\to|R|$") measures the average strength of the relation between current and future absolute returns (a proxy for volatility). The leverage effect expects this relation to be negative: a large price drop (negative return) is followed by periods of increased volatility (absolute returns). It also expects the average $L(k)$ for $k>0$ (labeled "$|R|\to R$") to be close to zero, indicating that current absolute returns do not forecast future returns.
@@ 4.5. Time Reversal Asymmetry
Time reversal asymmetry (TRA) in financial time series, introduced by [21], is a concept linked to entropy and time directionality. It suggests that financial time series exhibit different statistical features when viewed in reverse chronological order. This phenomenon can be illustrated by the analogy of observing broken glass: while one can easily deduce that glass fragments originated from an intact cup, it’s nearly impossible to envision the original cup when presented only with its fragments.
@@
To quantify TRA, we employ a correlation measure introduced in [13]:
$C(k)\equiv\frac{\mathbb{E}\left[(\lvert R_{t}\rvert-\mu_{\lvert R\rvert})(s^{(D)}_{t+k}-\mu_{s^{(D)}})\right]}{\sigma_{\lvert R\rvert}\sigma_{s^{(D)}}}$ (3)
@@
An increasing function of $\Delta(N)$ confirms the presence of TRA [13].
Figure 11 illustrates $\Delta(N)$ for selected assets over the same historical period. The consistently increasing lines for all assets, including both crypto and traditional financial instruments (SPY and EUR), validate the presence of TRA, aligning with previous findings in traditional finance [21, 13].
Figure 11. Cumulative difference function $\Delta(N)$ (Equation (4)) for ETH in different exchanges as well as SPY and EUR.
@@ 4.6. Correlation Matrix and Market Factors
Figure 14 reveals striking differences between the token and stock markets. In the token universe, the top eigenvalue accounts for 56% of the total variance, whereas in the stock samples, it explains only about 26% (reported conservatively as the average minus three standard errors over the 500 randomly selected stock portfolios). Importantly, the first eigenvector is positive for all tokens in our sample, consistent with its interpretation as the "market" factor, as discussed in [14, 5, 3].
@@
These findings suggest that the token universe is driven by a smaller set of eigenvalues/factors than the stock market while still adhering to the stylized fact of factor-driven markets observed in stocks [14, 5, 3]. This conclusion has important implications for modeling and risk control of token portfolios. Specifically, it suggests that a few factor models—likely fewer than those required for stocks—may be sufficient to effectively model the cross-section of tokens.
4.6.1. Correlation Across Time
While illustrating the temporal evolution of the correlation matrix presents challenges due to limited reliable data, we can gain qualitative insights by analyzing a subset of our dataset. We focus on 25 tokens with data beginning in September 2020, extending our analysis two years earlier than the data used in Figure 14. This period encompasses significant market phases, including a bull market that saw Bitcoin’s price increase nearly five-fold, followed by substantial volatility in early 2021, and eventually establishing a bull market toward the end of 2021.
To analyze correlation dynamics across these market regimes, we examine the time series of the first eigenvalue of the correlation matrix for our 25-token subset. Our methodology involves calculating a 60-day trailing correlation matrix daily, computing its eigenvalue spectrum, and recording the fraction of variance explained by the first eigenvalue at the end of each 60-day period.
Figure 15. Time-series of the variance explained by the first eigenvalue for a rolling 60-day correlation matrix of 25 tokens (top) with the cumulative sum of the log-returns for the same portfolio of 25 tokens (lower). The graph illustrates that bull markets are more diversified (less variance explained by the first eigenvalue) and bear markets are less diversified (more variance explained by the first eigenvalue).
Figure 15 presents two key elements: the time series of the fraction of variance explained by the first eigenvalue (upper panel) and the cumulative sum of log-returns averaged across our 25-token sample (lower panel). During the bear period after 2022, the average explained variance approaches 66%, notably higher than the 56% observed in our previous section. This difference stems from our use of a substantially smaller, less diversified correlation matrix in this analysis.
Despite the limited data span, our analysis reveals distinct patterns in market behavior. The first eigenvalue tends to be smaller (approximately 50%) during bull markets and higher (around 65%) during bear markets. This transition became particularly evident during the significant market downturn in 2021. Subsequently, the fraction of explained variance has fluctuated around 65%, occasionally reaching 80% during rapid market declines. These findings demonstrate a negative relationship between token market performance (particularly Bitcoin) and the fraction of explained variance. During substantial market upswings, tokens show more independent movement, while market crashes tend to trigger synchronized downward movements across tokens.
This breakdown in diversification benefits during market crashes mirrors well-documented patterns in traditional stock markets [15]. Our analysis suggests that Web3 tokens exhibit similar behavior, where diversification benefits diminish precisely when they would be most valuable – during market downturns.